OpenStudy (ambermarie151):

If you shift the quadratic parent function, f(x) = x2, right 10 units, what is the equation of the new function? A. g(x) = x^2 + 10 B. g(x) = (x – 10)^2 C. g(x) = x^2 – 10 D. g(x) = (x + 10)^2 I think it's B

OpenStudy (danjs):

A horizontal shift in the x direction for a function y = f(x), will change the x values by the amount of shift 'b' f(x - b) = (x - 10)^2

OpenStudy (ambermarie151):

So it will change to a positive ? @DanJs

OpenStudy (danjs):

It is the oposite sign you think, to go right you take f(x - 10), to go left you take f(x - (-10)) = f(x + 10)

OpenStudy (jackthegreatest):

the equation is x-h, where h is the shift, so if it is to the right, it will appear negative

OpenStudy (ambermarie151):

I am so confused @jackthegreatest Danjs

OpenStudy (danjs):

if you slide the whole graph horizontally in the x direction, all the current X values will be moved to (x - b) instead. Where the shift is 'b' units in the horizontal direction.

OpenStudy (danjs):